OpenStudy (princesssleelee):

4 times the square root of the quantity of x plus 2 equals negative 16,

OpenStudy (princesssleelee):

\[4\sqrt{x+2}=-16\]

OpenStudy (princesssleelee):

@SolomonZelman

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

do u have to simplify it ?

OpenStudy (princesssleelee):

find the solution

OpenStudy (anonymous):

are there any answer choices?

OpenStudy (princesssleelee):

x = 14; not extraneous x = 2; extraneous x = 2; not extraneous x = 14; extraneous

OpenStudy (anonymous):

oh u want to solve for x

OpenStudy (anonymous):

its 14 not extraneous

OpenStudy (princesssleelee):

how do you figure

OpenStudy (anonymous):

you divide 4 from both sides, then square both sides, then subtract 2 from both sides

OpenStudy (princesssleelee):

i was thinking it was ext.

OpenStudy (anonymous):

ok i ll go step by step with u ok bear with me

OpenStudy (anonymous):

first square both side so u dont have to deal with sqrt \[16(x+2)=256\]

OpenStudy (anonymous):

sorry it is extraneous because when you plug 14 into the equation it doesnt equal -16 it equals 16

OpenStudy (anonymous):

divide both sides by 16 \[x+2=\frac{ 256 }{ 16 }\]

OpenStudy (anonymous):

then \[x+2=16\]

OpenStudy (anonymous):

\[x=16-2\]

OpenStudy (princesssleelee):

D looks good to me, but continue @ayeshaafzal221 , lets see if you concluded the same answer :)

OpenStudy (anonymous):

x=14

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

its not extraneous we dont have two solution

OpenStudy (anonymous):

as u saw from my steps we only got 14

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

No its A

OpenStudy (anonymous):

dude plug it in to the equation

OpenStudy (anonymous):

its extraneous

OpenStudy (princesssleelee):

lol i got D as well

OpenStudy (anonymous):

D is correct

OpenStudy (anonymous):

no because i squared it

OpenStudy (princesssleelee):

i have a couple other questions if yall are possibly able to help :)

OpenStudy (anonymous):

I could try, what grade is this?

OpenStudy (anonymous):

oh wait @SchoolHelpPls is right srry about that i was thinking of something else

OpenStudy (princesssleelee):

Simplify (5x3− x + 14) − (3x2− 9x + 4). 5x3− 3x2+ 8x + 10 2x2+ 8x + 10 5x3− 3x2− 10x + 18 2x2− 10x + 18

OpenStudy (anonymous):

C

OpenStudy (shadowlegendx):

@princesssleelee Put that into a new question. Itʻs not against rules, but it keeps things more organized @SchoolHelpPls No direct answers

OpenStudy (anonymous):

sorry A i misread

OpenStudy (shadowlegendx):

@SchoolHelpPls Explain your answers, even Wolfram doesnʻt explain...

OpenStudy (princesssleelee):

Its okay, I got A as well honestly i just needed a double check on one so i know that i did the others like it correctly :)

OpenStudy (anonymous):

@ShadowLegendX still gives the right answers tho ;) and in adulthood you only use 3% of what you learn in school so who cares. School only cares about your grades anyway

OpenStudy (princesssleelee):

Very true!

OpenStudy (princesssleelee):

9/3-2i

OpenStudy (princesssleelee):

i got 27+18i/13

OpenStudy (anonymous):

@SchoolHelpPls but she is not learning anything if u give her straight answers , next time she wont need to ask if she knows how to do it !

OpenStudy (shadowlegendx):

@SchoolHelpPls Where did you get that 3% statistic? Also, on her tests and exams, these direct answers wonʻt get her an A. Good luck on SATs / ACTs ._.

OpenStudy (princesssleelee):

I just passed my ACT!

OpenStudy (anonymous):

A lot of Collages don't even require you to take those tests anymore to enter because people are realizing that standardized tests are bs

OpenStudy (shadowlegendx):

@SchoolHelpPls Where are you getting this information

OpenStudy (princesssleelee):

So no more help for me? lol i posted a problem and the answer i got ..... lol

OpenStudy (shadowlegendx):

@princesssleelee What of upcoming test and exams? and at college you will need to know this stuff as well. Donʻt just blow it off, it will hurt you in the future, trust me

OpenStudy (princesssleelee):

ok..... soooooo i posted a problem and the answer that i got... could i be told if i got the correct answer or where i may have made an error?

OpenStudy (anonymous):

@princesssleelee I'm sorry, I'm not sure how to do the problem, I'm in 10th and haven't seen that yet =/

OpenStudy (anonymous):

can you type it on here exactly how it looks on the test?

OpenStudy (princesssleelee):

\[\frac{ 9 }{ 3-2i }\]

OpenStudy (princesssleelee):

i got \[\frac{ 27+18i }{ 13 }\]

OpenStudy (anonymous):

Thank you! do you have answer choices also?

OpenStudy (shadowlegendx):

a 10th grader thinks itʻs best to judge a person on what he/she should know for the future...

OpenStudy (anonymous):

You are correct @princesssleelee

OpenStudy (princesssleelee):

it cannot be simplified anymore, correct?

OpenStudy (anonymous):

I can't see how if it can be. I'm 99% sure the answer you have is correct

OpenStudy (princesssleelee):

EVERYTHING you gave me/worked with me on were perfect! i got one wrong on the entire thing..... a 97% but i did that one on my own thinking it was right, should of had you check it ;) lol thank you very much! it was greatly appreciated, i know how to do every problem because of you! even with the answer, i still worked out the problem the right way! and found where i was wrong when i didnt!

OpenStudy (anonymous):

You're very welcome! Good luck in the future and Happy Holidays! :)

OpenStudy (princesssleelee):

To you as well! I hope youre on often, however i have until january 26th to finish this class lol!